# Sequential model confidence sets

### v1

@@ 4.2 Simulation 2
We consider the simulation setting as given in Example 2.4. By (4.1),
$\mathbb{E}(L_{i,t})$ $=\mathbb{E}\big{(}Z_{t}\Phi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}-\varepsilon_{i}\mathbb{E}\big{(}\Phi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}+\varepsilon_{i}+2\sqrt{1+\delta_{i}}\mathbb{E}\big{(}\varphi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}-\sqrt{\frac{1+\delta_{i}}{\pi}}$
$=\frac{1+2\sqrt{1+\delta_{i}}}{\sqrt{2\pi(2+\delta_{i})}}\exp\left\{-\frac{\alpha_{i}^{2}}{2}\right\}-\varepsilon_{i}\Phi\left(-\alpha_{i}\right)+\varepsilon_{i}-\sqrt{\frac{1+\delta_{i}}{\pi}},$ (16)
where $\alpha_{i}=\varepsilon_{i}/\sqrt{2+\delta_{i}}$ and $\Phi_{\mu,\sigma^{2}}$ ($\varphi_{\mu,\sigma^{2}}$) denotes the cdf (density) of $\mathcal{N}(\mu,\sigma^{2})$ for $\mu\in\mathbb{R}$ and $\sigma>0$. For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, let $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mu_{ij,t}$. Since the CRPS is a proper scoring rule, we have $\mu_{i_{0}j,t}\leq 0$ for all $t\notin 7\mathbb{N}$, for all $j\in\mathcal{M}_{0}$. By (16), for $j\in\mathcal{M}_{0}$ with $(\varepsilon_{j},\delta_{j})\in\{-0.2,0.2\}^{2}$, we have $\mu_{i_{0}j,t}>0$ for all $t\in 7\mathbb{N}$. Thus, $\mathcal{M}^{\textrm{s},\star}_{t}=\emptyset$ for $t\geq 7$. However, equation (16) allows to show numerically that $\Delta_{i_{0}j,t}\leq 0$ for all $j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, and thus $\mathcal{M}^{\textrm{uw},\star}_{t}=\{i_{0}\}$ for all $t\in\mathbb{N}$.
We compute SMCSs for $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\leq n}$ as suggested in Section 3.3 with confidence level $\alpha=0.1$. We transform the conditionally bounded CRPS differences (see Appendix B) as proposed in Section 3.5 into uniformly bounded score differences $\lvert\tilde{d}_{ij,t}\rvert\leq C/2$ for $C=2$, and let $\lambda=(2C)^{-1}=1/4$. Figure 1 displays the average size of the SMCS for $N=100$ repetitions of the simulation. Again, we have a 100% coverage rate of the superior model $i_{0}$. Notably, we have less power compared to the SMCS for the strongly superior objects and still remain with approximately $10$ surviving models after $1000$ observations.
@@ 4.3 Simulation 3
We consider the simulation setting as given in Example 2.5. Let $i=1,2,3$ be forecasters issuing median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, $t\in\mathbb{N}$, for i.i.d. standard normally distributed $(Y_{t})_{t\in\mathbb{N}}$, where the biases $\varepsilon_{i,t}$ are given in (6). For $t\in\mathbb{N}$, we have by independence
$2\mu_{ij,t}=\mathbb{E}(\Phi(m_{i,t}))-\mathbb{E}(\Phi(m_{j,t}))=\Phi\left(\frac{\varepsilon_{i,t}}{\sqrt{2}}\right)-\Phi\left(\frac{\varepsilon_{j,t}}{\sqrt{2}}\right),\quad i,j=1,2,3,$
and thus
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i=1,2,3\mid\sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{i,t}}{\sqrt{2}}\right)\leq\sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{j,t}}{\sqrt{2}}\right),j=1,2,3\right\}.$ (17)
We let $\beta=0.6,\gamma=0.998$ and $\delta=0.008$. Then, one can numerically show by (17) that $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$.
We sample $n=800$ observations and construct SMCSs for $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\leq n}$ at level $\alpha=0.1$ as suggested in Section 3.4. We use the fact that $(d_{ij,t})_{t\in\mathbb{N}}$ is uniformly bounded with $c=1$ and let $\lambda_{ij}=1.1^{-1}$, $i,j=1,2,3$. Figure 2 shows the accumulated observed losses and the resulting SMCS for one realization of the simulation (left), and the average size of the SMCS over time for $N=100$ runs (right). Again, the SMCS includes the theoretical superior objects given at (17) at all time points.
Refer to captionRefer to caption
Figure 2: For $i=1,2,3$, the left panel shows the realized accumulated losses $\sum_{r=1}^{t}L_{i,r}$, $t\leq n$, for one simulation. The black vertical lines indicate $t=154$ and $t=550$. The resulting SMCS is given in the upper part with the respective colors: The SMCS correctly excludes the improving forecaster (blue) at the beginning and the disimproving forecaster (green) relatively quickly after the first change point until it includes the improving forecaster again and finally excludes the constantly biased forecaster (red). The right panel shows the average size of the SMCS over $N=100$ realizations, where we again have a $100\%$ coverage rate of the superior models.
@@ 5.1 Covid-19 case study
After the outbreak of the Covid-19 pandemic, the Reich lab at the University of Massachusetts collaborated with the United States centers for disease control and prevention (CDC) to create the Covid-19 forecast hub, a repository containing point and probabilistic forecasts for incident cases, hospitalizations, and Covid-19 related deaths. The repository was founded in March 2020, and attracted the submission of forecasts by more than $60$ different teams. The collection of forecasts is still ongoing, although most of the models had stopped submitting forecasts by January 2024, due to the subsidence of the pandemic.
5.1.1 The data
@@
The dataset, which we use in this application, is discussed in detail in Cramer et al. (2022a) and Cramer et al. (2022b), and publicly available at the Covid-19 forecast hub GitHub repository and on the Zoltar forecast archive. It contains forecasts with corresponding observations for key epidemiology indices for 55 different locations in the US as well as on the aggregated US national level. The forecasts are issued by a total of 69 different forecast models and reported at $23$ different quantile levels with forecast horizons ranging from $1$ day to $20$ weeks. Importantly, many models are not directly comparable since not all of them have issued their forecasts on the same weekdays or throughout the whole period.
@@
In this case study, we consider 1-week-ahead forecasts for Covid related deaths on the national level, issued weekly in the period from 06/06/2020 to 04/03/2024, and focus on the comparison of the following $m=6$ different forecasting models: Firstly, we consider the benchmark model baseline, which naively issues the most recent outcome as the median prediction for the following week and forms a predictive distribution around this median prediction based on the past weekly incidences. Secondly, we consider two epidemic models PSI-draft and MOBS-gleam issuing their forecasts based on epidemiological assumptions, and the neural network based model GT-deep, which was the first purely data-driven model to be included in the Covid-19 forecast hub. Finally, we consider the summary models ensemble and CDC-ensemble, where the former combines the most recent predictions of all other models into one predictive distribution, and the latter is supposed to improve the former by only considering the 10 current best models, measured by the average weighted interval score over the 12 most recent weeks. For a more detailed documentation of the models, we refer to Appendix F.
@@
Regarding the nature of the models, we perform our comparison under the strong hypothesis, since we expect the ensemble models to have consistent superior performance over the individual models, which in turn are expected to be better than the naive baseline model issuing its forecasts only based on the previous outcomes and on no additional information. Furthermore, we can argue that in such a highly sensible prediction task, we should seek to detect the models that issue accurate predictions consistently over time and not just on average.
@@
Notably, the outcomes as well as the forecasts in the given dataset have a very unstable behaviour and exhibit large variation in scale, see e.g. the upper left panel of Figure 3, which suggests to apply a log transformation on all quantities of interest. Moreover, the forecast differences are typically not very distinctive for the median, whereas they become more apparent when we look at the tails, compare to the lower panels in Figure 3.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 3: Time progression of predicted and actual Covid related mortality in linear (upper left) and log scale (upper right). For the forecasts at the tails, the differences between the different models are more pronounced than for the median, see the lower panels for $\tau=0.15,0.975$.
@@ 5.1.2 Methods and implementation details
For $\tau\in(0,1)$, we assess the different quantile forecasts at level $\tau$ with respect to the generalized piecewise-linear (GPL) quantile scoring function (Gneiting, 2011),
$S_{\tau}(x,y)=\left(\mathbf{1}\{x\geq y\}-\tau\right)\left(\log(x)-\log(y)\right),\quad x,y\in\mathbb{R}.$
The log transformation may thus easily be justified from a statistical point of view by simply setting $g=\log$ in the formula of the GPL score, see Appendix A for details. In Appendix B, it is shown that the resulting quantile score differences are conditionally bounded. More precisely, for $\tau$-quantile predictions $(x_{i,t})_{t\in\mathbb{N}},(x_{j,t})_{t\in\mathbb{N}}\subseteq\mathbb{R}$ issued by the models $i,j=1,\dots,m$, we have $\lvert d_{ij,t}\rvert=\lvert S_{\tau}(x_{i,t},y_{t})-S_{\tau}(x_{j,t},y_{t})\rvert\leq c_{ij,t}/2$ for
@@
We apply our methods of Section 3.2 to construct SMCSs for the strongly superior objects, where we use the e-processes $E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r})$ for $(\lambda_{ij,t})_{t\in\mathbb{N}}$ predictable with $0<\lambda_{ij,t}<c_{ij,t}^{-1}$, $t\in\mathbb{N}$. For $i,j=1,\dots,m$ and $t\in\mathbb{N}$, we suggest to set
$\lambda_{ij,t}=\frac{1}{K_{ij,t}c_{ij,t}+\varepsilon}$
for some small $\varepsilon>0$, included as a safeguard for the scenario where the two forecaster issue the same prediction and thus $c_{ij,t}=0$, and for
@@
The coefficient $K_{ij,t}$ depends on the sign and magnitude of the most recent score difference $d_{ij,t-1}$ as well as of the centrality of $\tau$: If $d_{ij,t-1}>0$, then we have evidence against the null hypothesis that model $i$ is strictly superior than model $j$, and the quantity $K_{ij,t}$ becomes small. That is, the parameter $\lambda_{ij,t}$ tends to be larger, and we bet more against the null hypothesis. On the other hand, $K_{ij,t}$ becomes larger for $d_{ij,t-1}<0$, which leads to smaller values of the tuning parameters $\lambda_{ij,t}$. Finally, the primary multiplicative factor, which takes values in $[1,2]$, is used to reduce the influence of the term $\max\left\{\tau,1-\tau\right\}$ on the bound in (18). It converges to $1$ as $\tau\downarrow 0$ or $\tau\uparrow 1$, and to $2$ as $\tau\to 0.5$. This means that we bet more at the tails where the forecasts are typically more distinctive and less if $\tau$ is close to $0.5$ where the forecasts are typically less distinguishable.
Choosing a smart betting strategy has an effect on the power of the methods as discussed in Section 3.5. In Appendix F, we compare the suggested choice of the predictable $\lambda_{ij,t}$’s with the naive approach where $K_{ij,t}=2$ for all $t\in\mathbb{N}$.
@@ 5.1.3 Results
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 4: The resulting SMCSs at four selected quantile levels with confidence level $\alpha=0.1$. Even though the SMCS behave differently at the different quantile levels, we may conclude that the ensemble models performed best as they are never excluded from the SMCSs. Second best is GT-deep, which only shows some deficiencies for $\tau=0.3$, and MOBS-gleam and the baseline, which seem to be approximately equally accurate. PSI-draft performs poorest and is excluded rather quickly at each quantile level.
@@
Figure 4 shows the resulting SMCSs for four selected quantile levels with a confidence level of $\alpha=0.1$. As a first observation, we see that the power of the SMCS varies with the choice of the quantile level $\tau$: Whereas the SMCS detects four deficient models for $\tau=0.3$, it only has excluded two models by the end of the observation period for $\tau=0.7$. For $\tau=0.1$, the SMCS needs considerably more observations to reject the three models, which are excluded much faster at the median. As argued before, the lower power at the tails of the distribution is probably due to the factor $\max\left\{\tau,1-\tau\right\}$ in the bound (18). In Appendix F, we show that the effect of lower power at the tails even increases if we do not incorporate the quantile level in our betting scheme and that our proposed betting scheme is capable to mitigate this difficulty. Overall, the results align with our expectation that the two ensemble models perform best as they are included in the SMCS for the entire time period at all quantile levels. The superior performance of the two ensemble models was expected since they combine the forecasts of all other models and thus have access to the largest information set. Interestingly, the SMCS does not detect any significant difference between ensemble and CDC-ensemble, where the latter was supposed to improve the former by combining just the currently best, rather than all, forecasting models. The naive baseline performs surprisingly well and even outperforms the epidemic model PSI-draft, which is excluded earlier than baseline for all given values of $\tau$. We may conclude that PSI-draft is not competitive with the other models and should therefore not be considered as an accurate forecasting model for the pandemic. Finally, the second epidemic model MOBS-gleam performs better than PSI-draft but not as good as the neural network based GT-deep according to the SMCSs.
@@
To conclude, we again highlight the key feature that SMCSs allow us to collect statistical evidence sequentially over time. Hence, if we were back in the pandemic, by using SMCSs, we could monitor the models performances from the very beginning on and would not have to wait until we have collected enough observations to perform a statistical test for predictive ability. Quickly, we would reject PSI-draft as an accurate model and not consider it in our analysis and policy-making anymore. Also MOBS-gleam would probably have been rejected as a competitive model by approximately the end of 2021, and so we would have been left with the choice between less and less models in the course of the pandemic. Clearly, in this extreme situation, it has already been common practice to assess the forecasting models on a regular basis. However, this practice may lack the theoretical justification, and our methodology provides statistically safe methods to do so.

### v3

@@ 4.2 Simulation 2
We consider the simulation setting as given in Section 2.1 for $n=1000$. By (11),
$\mathbb{E}(L_{i,t})$ $=\mathbb{E}\left(Z_{t}\Phi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)\right)-\varepsilon_{i}\mathbb{E}\Phi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)+\varepsilon_{i}+2\mathbb{E}\varphi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-\sqrt{\frac{1+\delta_{i}}{\pi}}$
$=\frac{1+2\sqrt{1+\delta_{i}}}{\sqrt{2\pi(2+\delta_{i})}}\exp\left\{-\frac{\alpha_{i}^{2}}{2}\right\}-\varepsilon_{i}\Phi\left(-\alpha_{i}\right)+\varepsilon_{i}-\sqrt{\frac{1+\delta_{i}}{\pi}}.$ (12)
For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, let $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mu_{ij,t}$. Since the CRPS is a proper scoring rule, we have $\mu_{i_{0}j,t}\leq 0$ for all $t\notin 7\mathbb{N}$ and $j\in\mathcal{M}_{0}$. By (12), for $j\in\mathcal{M}_{0}$ with $(\varepsilon_{j},\delta_{j})\in\{-0.2,0.2\}^{2}$, we have $\mu_{i_{0}j,t}>0$ for all $t\in 7\mathbb{N}$. Thus, $\mathcal{M}^{\textrm{s},\star}=\emptyset$. However, equation (12) allows to show numerically that $\Delta_{i_{0}j,t}\leq 0$ for all $j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, and thus $\mathcal{M}^{\textrm{uw},\star}=\{i_{0}\}$.
We compute SMCSs for $\mathcal{M}^{\textrm{uw},\star}$ with confidence level $\alpha=0.1$ using the approach in Section 3.2. We transform the conditionally bounded CRPS differences (see Supplement B) as proposed in Section 3.4 into uniformly bounded score differences $\lvert\tilde{d}_{ij,t}\rvert\leq c/2$ for $c=2$, and let $\lambda_{ij}=(2c)^{-1}=1/4$. Figure 1 displays the average SMCS size across time, which declines slightly slower than in Simulation 1. This is to be expected, since we work with a weaker hypothesis. Regarding coverage of the superior object, the SMCS is still conservative and exhibits a coverage rate of $100\%$.
It could be interesting to monitor the e-processes that we use to collect evidence on predictive performance over time. Due to the large number of models considered here, this is not so practical, but we illustrate this idea in Supplement F.1, where only nine models are considered in the experiment.
@@ 4.3 Simulation 3
We consider the simulation setting from Section 2.1. Let $i=1,2,3$ be forecasters issuing median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, $t\in\mathbb{N}$, for i.i.d. standard normally distributed $(Y_{t})_{t\in\mathbb{N}}$ and $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $10<\gamma<$ 1. For the loss $L(m,y)=0.5(\Phi(m)-\Phi(y))$, we get by independence that $2\mu_{ij,t}=\mathbb{E}(\Phi(m_{i,t}))-\mathbb{E}(\Phi(m_{j,t}))=\Phi\left(\varepsilon_{i,t}/\sqrt{2}\right)-\Phi\left(\varepsilon_{j,t}/\sqrt{2}\right)$ for $t\in\mathbb{N}$, $i,j=1,2,3$, thus
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i\ \middle|\ \sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{i,t}}{\sqrt{2}}\right)\leq\sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{j,t}}{\sqrt{2}}\right),\;\text{for all}\;j\right\}.$ (13)
We let $\beta=0.6,\gamma=0.998$ and $\delta=0.008$. Then, one can numerically show that $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$.
We sample $n=800$ observations and construct the SMCSs defined in Section 3.3 for $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\leq n}$, at level $\alpha=0.1$. We use the fact that $(d_{ij,t})_{t\in\mathbb{N}}$ is uniformly bounded with $c=1$ and let $\lambda_{ij}=1.1^{-1}$, $i,j=1,2,3$. Figure 2 shows the accumulated observed losses $\sum_{r=1}^{t}L_{i,r}$ and the resulting SMCS for one realization. The SMCS correctly excludes the improving forecaster at the beginning and the worsening forecaster relatively quickly after the first change point ($t=154$) until it includes the improving forecaster again and finally excludes the constantly biased forecaster. The average size of the SMCS for $N=100$ runs varies between one and two models after a short initial period with three models. The SMCS includes the superior objects given at (13) at all time points.
Figure 2: Left: Realized accumulated losses $\sum_{r=1}^{t}L_{i,r}$ for one realization: Worsening forecaster (green), improving forecaster (blue), constantly biased forecaster (red). The black vertical lines indicate $t=154$ and $t=550$. The resulting SMCS is given in the upper part with the respective colors. Right: Average size of the SMCS over $N=100$ realizations.
@@ 5.1 Covid-19 case study
After the outbreak of the Covid-19 pandemic, the Reich lab at the University of Massachusetts collaborated with the United States centers for disease control and prevention (CDC) to create the Covid-19 forecast hub, a repository containing point and probabilistic forecasts for incident cases, hospitalizations, and Covid-19 related deaths. The repository was founded in March 2020, and attracted submissions of forecasts from more than $60$ different teams. The collection of forecasts is still ongoing, although most of the models had stopped submitting forecasts by January 2024, due to the subsidence of the pandemic.
5.1.1 Data
@@
The dataset we use in this application, is discussed in detail in Cramer et al. (2022a, b), and publicly available at the Covid-19 forecast hub GitHub repository and on the Zoltar forecast archive. It contains forecasts with corresponding observations for key epidemiology indices for 55 different locations in the U.S. as well as on the aggregated U.S. national level. The forecasts are issued by a total of 69 different forecast models and reported at $23$ different quantile levels with forecast horizons ranging from $1$ day to $20$ weeks. Some models are not statistically comparable since they have consistently issued their forecasts on different weekdays. There are also models that started issuing forecasts late into the forecasting period and others that stop early. Finally, there are models with a large number of missing values. In principle, one could use missing value imputation, or assign a default score to missing values so that the number of comparable models is higher. However, the purpose of this study to demonstrate how sequential model confidence sets work, so we so not go further in this direction. We only include models that are directly comparable without any preprocessing.
@@
Specifically, we consider 1-week-ahead forecasts for Covid related deaths on the national level, issued weekly in the period from 06/06/2020 to 04/03/2024, and focus on the comparison of the following $m=6$ different forecasting models: Firstly, we consider the benchmark model baseline, which naively issues the most recent outcome as the median prediction for the following week and forms a predictive distribution around this median prediction based on the past weekly incidences. Secondly, we consider two epidemic models PSI-draft and MOBS-gleam issuing their forecasts based on epidemiological assumptions, and the neural network based model GT-deep, which was the first purely data-driven model to be included in the Covid-19 forecast hub. Finally, we consider the summary models ensemble and CDC-ensemble, where the former combines the most recent predictions of all other models into one predictive distribution, and the latter is supposed to improve the former by only considering the 10 current best models, measured by the average weighted interval score over the 12 most recent weeks. For a more detailed documentation of the models, we refer to Cramer et al. (2022a).
@@
We perform our comparison under the strong hypothesis, since we expect the ensemble models to have consistent superior performance over the individual models, which in turn are expected to be better than the naive baseline model issuing its forecasts only based on the previous outcomes and on no additional information. Furthermore, we can argue that in such a highly important prediction task, we should seek to detect the models that issue accurate predictions consistently over time and not just on average.
@@
The outcomes as well as the forecasts exhibit unstable behaviour and large variation in scale, see e.g. the upper left panel of Figure 3, which suggests to apply a log transformation on all quantities of interest. Moreover, forecast differences are typically not distinctive for the median, while they become more apparent when we look at the tails, see the lower panels in Figure 3.
@@
Figure 3: Time progression of predicted and actual Covid related mortality in linear (upper left) and log scale (upper right). For the forecasts of the tails, the differences between the different models are more pronounced than for the median, see the lower panels for $\tau=0.15,0.975$.
@@ 5.1.2 Methods and implementation details
We assess the different quantile forecasts at level $\tau\in(0,1)$ with respect to the generalized piecewise-linear quantile scoring function $S_{\tau}(x,y)=\left(\mathbf{1}\{x\geq y\}-\tau\right)\left(\log(x)-\log(y)\right)$ for $x,y\in\mathbb{R}$, see Gneiting (2011) and Supplement A. In Supplement B, it is shown that the resulting quantile score differences are conditionally bounded. More precisely, for $\tau$-quantile predictions $(x_{i,t})_{t\in\mathbb{N}},(x_{j,t})_{t\in\mathbb{N}}\subseteq\mathbb{R}$ issued by the models $i,j=1,\dots,m$, we have $\lvert d_{ij,t}\rvert=\lvert S_{\tau}(x_{i,t},y_{t})-S_{\tau}(x_{j,t},y_{t})\rvert\leq c_{ij,t}/2$ for
@@
We apply our methods from Section 3.2 to construct SMCSs for the strongly superior objects, using the e-processes $E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r})$ for a predictable sequence $(\lambda_{ij,t})_{t\in\mathbb{N}}$ with $0<\lambda_{ij,t}<c_{ij,t}^{-1}$, $t\in\mathbb{N}$. For $i,j=1,\dots,m$ and $t\in\mathbb{N}$, we suggest $\lambda_{ij,t}=({K_{ij,t}c_{ij,t}+\varepsilon})^{-1}$ for some small $\varepsilon>0$, included as a safeguard for the scenario $c_{ij,t}=0$, which would be the case if the two forecasters issued the same prediction, and
@@
The coefficient $K_{ij,t}$ depends on the sign and magnitude of the most recent score difference $d_{ij,t-1}$ as well as of the centrality of $\tau$: If $d_{ij,t-1}>0$, then we have evidence against the null hypothesis that model $i$ is strictly superior than model $j$, and the quantity $K_{ij,t}$ becomes small. That is, the parameter $\lambda_{ij,t}$ tends to be larger, and we bet more against the null hypothesis. Finally, the primary multiplicative factor, which takes values in $[1,2]$, is used to reduce the influence of the term $\max\left\{\tau,1-\tau\right\}$ on the bound in (14).
@@ 5.1.3 Results
Figure 4: SMCSs at four selected quantile levels with confidence level $\alpha=0.1$.
@@
Figure 4 shows the resulting SMCSs for four selected quantile levels with a confidence level of $\alpha=0.1$. As a first observation, we see that the power of the SMCS varies with the choice of the quantile level $\tau$: Whereas the SMCS detects three deficient models for $\tau=0.1,0.3,0.5$, it only has excluded two models by the end of the observation period for $\tau=0.7$. For $\tau=0.1$, the SMCS needs considerably more observations to reject the three models, which are excluded much faster at the median. The lower power in the tails of the distribution may be due to the factor $\max\left\{\tau,1-\tau\right\}$ in the bound (14). Overall, the results show that the two ensemble models perform best, as they are included in the SMCS across the entire time period, and at all quantile levels. The superior performance of the two ensemble models was expected, since they combine the forecasts of all other models and thus have access to a larger information set. Interestingly, the SMCS does not detect any significant difference between ensemble and CDC-ensemble, where the latter was supposed to improve the former by combining only the best-performing forecasting models. The naive baseline performs surprisingly well and even outperforms the epidemiological model, PSI-draft, which is excluded earlier than the baseline for all given values of $\tau$. We may conclude that PSI-draft is not competitive against the other models and should therefore not be considered as an accurate forecasting model for the pandemic. Finally, the second epidemiological model, MOBS-gleam, performs almost always better than PSI-draft but not as good as the neural network based GT-deep.
@@
To conclude, we highlight once again the key feature that SMCSs allow us to collect statistical evidence sequentially over time. Hence, if we were back in the pandemic, by using SMCSs, we could monitor the models performances from the beginning on and would not have to wait until we have collected enough observations to perform a statistical test for predictive ability. Quickly, we would reject PSI-draft as an accurate model and not consider it in our analysis and policy-making anymore. Also MOBS-gleam would probably have been rejected as a competitive model by approximately the end of 2021. Clearly, in such emergency situations, it has already been common practice to assess the forecasting models on a regular basis. However, this practice may lack theoretical justification, and our methodology provides statistically safe methods to do so.
We focused primarily on the sequential monitoring of model performances but our methodology also allows for optional stopping. For example, a possible stopping criterion back in the pandemic could have been the first day when there are only half of the original models left.
Throughout the paper, we have used a significance level equal to $\alpha=0.1$, but clearly, there is freedom to do otherwise. In particular, we could hope to detect some defective models even earlier by increasing $\alpha$ if we are willing to increase the type-I error control level.